OpenStudy (anonymous):

The are n teams in a league and the number of games played is given by the polynomial n^2-n=N. Women's softball league has 13 teams. what is the total of games played.

OpenStudy (anonymous):

N=(13)^2-13=156 gamee

OpenStudy (blacksteel):

Well, you just have to plug n into the equation: N = 13^2 - 13 = 156.

OpenStudy (anonymous):

thank you